hello, i dont know how to solve this question
x^3/4=125
please help =(

Found 2 solutions by nerdybill, Alan3354:
Answer by nerdybill(7384)

x^3/4=125
Raised both sides to the fourth power:
(x^3/4)^4 = (125)^4
x^3 = (125)^4
Raising both sides to the 1/3 power:
(x^3)^1/3 = ((125)^4)^1/3
x = 125^4/3
x = 125(125^1/3)
x = 125((5^3)^1/3)
x = 125(5)
x = 625

Answer by Alan3354(69443)

x^3/4=125
x^3/4=5^3
x^1/4=5
x = 5^4
x = 625